a chemist has 40 ml of a solution that is 50% chlorine. How much water should he add to make a solution that is 10% chlorine?
Answer by checkley77(12844)

.5*40=.10(40+x)
20=4+.10x
.10x=20-4
.10x=16
x=16/.10
x=160 ml of water is used.
Proof.
.5*40=.10(40+160)
20=.10*200
20=20